OpenStudy (anonymous):

Solve by the multiplication with add or subtract method 3x - 4y = 45 7x + 6y = 13

OpenStudy (wolf1728):

A) 3x - 4y = 45 B) 7x + 6y = 13 Multiply A) by 1.5 A) 4.5x -6y = 67.50 B) 7x +6y =13 Ading both equations 11.5x = 80.50 x = 7 Substituting 7 into equation B 49 + 6y = 13 6y = 36 y = -6

OpenStudy (anonymous):

Thank you guys so much

OpenStudy (j2lie):

Is this a substitution or elimination problem?

OpenStudy (anonymous):

Just the multiplication with adding and substracting

OpenStudy (j2lie):

oh ok.